Calculate the probabilities of P(z>z0) = 0.7389
1
Expert's answer
2021-10-01T08:04:31-0400

"P(Z>Z_0)=0.7389\\

\implies 1-P(Z≤Z_0)=0.7389\\

\implies P(Z≤Z_0)=1-0.7389=0.2611\\

\implies Φ(Z_0)=0.2611\\

\implies Z_0=Φ^{-1}(0.2611)=-0.639958"


Answer "Z_0= - 0.64"